After reading the​ Novy-Marz and Rauh paper regarding unfunded pension liabilities you decide to compute the total obligation of the state that you live in. After some research you determine that your​ state's promised pension payments amount to $1.5 billion dollars​ annually, and you expect this obligation to grow at 2.1% per year. You determine that the riskiness of this obligation is the same as the riskiness of the​ state's debt. Based on the pricing of that debt you determine that the correct discount rate for the​ fund's liabilities is 3.2% per annum.​ Currently, based on actuarial calculations using 8% as the discount​ rate, the plan is neither over nor underfundeddash–the value of the liabilities exactly matches the value of the assets. What is the extent of the true unfunded​ liability?
Answer #1

Value of the liability actually = First year liability / (r - g) = 1.5 / (3.2% - 2.1%) = $ 136.36 billion

Value of the funded liability = value of the liability using discount rate of 8% = 1.5 / (8% - 2.1%) = $ 25.42 billion

Hence, the extent of the true unfunded​ liability = $ 136.36 - 25.42 = $ 110.94 billion